# Order-agnostic Identifier for Large Language Model-based Generative Recommendation

### v1

@@ 1. Introduction
Token-sequence identifier utilizes a discrete token sequence to represent multi-dimensional item information. To generate items, LLMs use beam search to generate the top- item identifiers. Despite the effectiveness, token-sequence identifiers suffer from the 1) local optima issue (Zhou and Hansen, 2005) in the beam search. As illustrated in Figure 1, beam search greedily selects the sequence with top- probabilities at each generation step. However, the initial tokens of the target identifier might not necessarily align with the user preference. As such, the prefix of the target identifier has a low probability and will be pruned by beam search, causing inaccurate results. 2) Low generation efficiency in the autoregressive generation, which requires multiple serial LLM calls, thereby causing unaffordable computing burdens (Lin et al., 2024e) and severely hindering real-world deployments.
@@
To this end, we propose SETRec, an effective implementation of the set identifier paradigm. To integrate semantic and CF information, SETRec leverages CF and semantic tokenizers to assign each item with an order-agnostic token set containing CF and semantic embeddings. To eliminate token dependencies, 1) for user history encoding, we propose a special sparse attention mask, which discards the visibility of other tokens within identifiers and retains access to previous identifiers. 2) For simultaneous token generation, we introduce a query-guided generation mechanism, which adopts learnable vectors to guide LLMs to generate the embedding for each specific information dimension. 3) To ground the generated embedding set to existing items, SETRec collects embeddings from all items as grounding heads to obtain the item scores for ranking. We instantiate SETRec on T5 and Qwen and evaluate it on four real-world datasets under various scenarios (e.g., full ranking, warm- and cold-start ranking, and diverse item popularity groups) to demonstrate the effectiveness, efficiency, and generalization ability. Additionally, we evaluate SETRec on Qwen with different model sizes (i.e., 1.5B, 3B, and 7B), exhibiting promising scalability on cold-start items as model size increases.
@@ 2. Preliminaries
where is the LLM vocabulary. Despite the effectiveness, generating token sequences would result in the local optima issue and inference inefficiency. As shown in Figure 3, continuously increasing the beam size slightly improves recommendation accuracy, but remains inferior to globally optimal results. Worse still, the token-by-token generation requires multiple serial LLM calls, which significantly lowers the inference speed and hinders real-world applications.
@@ 4. Experiment
We carry out extensive experiments on four real-world datasets to answer the following research questions:
RQ1: How does our proposed SETRec perform compared to different identifier baselines on different architectures of LLMs?
RQ2: How do the different components of SETRec (i.e., CF embeddings, semantic embeddings, query vectors, and sparse attention) affect the performance?
RQ3: How does SETRec perform when scaling up the model size and how does SETRec improve the overall performance?
RQ4: How does SETRec perform with different number of semantic embeddings, tokenizer training strength, and semantic strength for inference?
@@ 4.1.1. Datasets
We conduct experiments on four real-world datasets across various domains. From Amazon review datasets222https://jmcauley.ucsd.edu/data/amazon/., we adopt three widely used benchmarks 1)Toys, 2) Beauty, and 3) Sports. The three Amazon datasets contain rich user interactions over a specific category of e-commerce products, where each item is associated with rich textual meta information such as title, description, category, and brand. In addition, we use a video games dataset 4) Steam333https://github.com/kang205/SASRec. proposed in (Kang and McAuley, 2018), which contains substantial user interactions on video games with abundant textual semantic information. For all datasets, we follow previous work (Wang et al., 2023) to sort user interactions chronologically according to the timestamps and divide them into training, validation, and testing sets with a ratio of 8:1:1. In addition, we divide the items into warm and cold items444We denote warm- and cold-start items as warm and cold items for brevity., where the items that appear in the training set are warm items, otherwise cold items.
@@ 4.3.2. Item Group Analysis (RQ3)
To understand how SETRec improves performance, we evaluate it over items with different popularity. We divide the items into 5 groups according to their frequencies and test the models over each group respectively. The performance comparison between SETRec and two competitive baselines from token-sequence identifiers (LETTER) and single-token identifiers (E4SRec) are reported in Figure 7. We can observe that 1) the performance gradually drops from G1 to G5. This makes sense since the less popular items have fewer interactions for LLMs to learn, thus leading to worse generation probabilities. Besides, 2) E4SRec outperforms LETTER on most popular items (G1) but usually yields inferior performance on unpopular items (G2-G5). This is due to that E4SRec only uses CF information, which relies on substantial interactions and therefore struggle on unpopular items. In contrast, LETTER additionally incorporates semantics into identifiers, thus achieving better generalization on sparse items. 3) SETRec consistently excels both E4SRec and LETTER over all groups. Notably, the improvements over sparse items are more significant, which partially explains the superiority of SETRec regarding overall performance.
@@ 4.3.4. Effect of Semantic Strength (RQ4)
To investigate how semantic information contributes to the performance during inference, we vary from to , where indicates that only CF score is used for ranking, and ranks items based solely on semantic scores (Eq. (7)). From the results reported in Figure 9. we can find that 1) Incorporating semantic information during inference is necessary (inferior performance of than , which facilitates global ranking over multi-dimensional information and lead to strong generalization ability. Notably, 2) incorporating semantic scores brings more significant improvements on cold items, underscoring the critical role of semantic information for zero-shot scenarios. Moreover, 3) Gradually increase to rely solely on semantics (), SETRec maintains competitive performance on warm items, which is probably attributed to the implicit alignment between CF and semantic tokens during training.
@@ 4.3.5. Hyper-parameter Sensitivity (RQ4)
We further study the hyper-parameter sensitivity to facilitate SETRec application.
Effect of . We vary the strength of AE loss for SETRec training and present the results on Toys in Figure 8(a-b). We can observe that 1) the performance is overall improved when is increased from to , which validates the effectiveness of reconstruction loss that encourages AE to preserve useful information in the latent space. Nonetheless, 2) while continuously increasing generally gives better performance on cold-start items, it might hurt the performance under “warm” setting. Based on the empirical results, we recommend setting ranging from to .
Effect of . We change the number of semantic tokens from to to investigate how affects the performance. From the results shown in Figure 8(c-d), we can find that 1) gradually increasing semantic tokens generally improves the performance, which validates the effectiveness of incorporating multiple tokens to mitigate the potential information conflicts (Wang et al., 2024a) and embedding collapse issue (Guo et al., [n. d.]). However, 2) blindly increasing the number of semantic tokens might hurt the performance (decreased performance from to ). This is reasonable since it is non-trivial to recover the category-level preference aligning well with the real-world scenarios. Similar observations are also seen in (Lin et al., 2024b) and (Lin et al., 2024d).
@@ 5. Related Work
LLM-based Recommendation. Harnessing LLMs for recommendations has garnered substantial attention across both academia and industry (Sun et al., 2024; Zhang et al., 2024c; Fu et al., 2024; Zhang et al., 2024c; Shi et al., 2024; Zhao et al., 2024; Zheng et al., 2024a; Li et al., 2024). Existing studies on LLM-based recommendations can be grouped into two research lines. 1) LLMs for discriminative recommendation, which typically aims to leverage LLMs to assist the conventional recommender models (Wang et al., 2024c; Ren et al., 2024a; Chen et al., 2024b; Wu et al., 2024). Harnessing LLMs’ strong reasoning ability, this line of work usually involves LLMs in different steps of recommendation pipelines (Lin et al., 2024a; Zhang et al., 2024a, b; Wang et al., 2024b) such as feature engineering (Xi et al., 2024; Ren et al., 2024b; Xi et al., 2024) and feature encoder (Chen et al., 2024a). 2) LLMs for generative recommendation, which regards LLMs as recommenders to directly generate items (Liao et al., 2024; Kim et al., 2024; Lin et al., 2024c; Wang et al., 2024f; Li et al., 2024). To build LLMs for generative recommendation, a key step is item tokenization, where each item is assigned an identifier for LLMs to encode user history and generate the next item. In this work, we critically analyze the fundamental principles of identifier design to achieve effective and efficient LLM-based generative recommendation.
@@
Identifier for LLM-based Recommendation. Existing identifier designs can be broadly categorized into two types: 1) token-sequence identifiers represent each item with a discrete token sequence. Under this group of work, prior effort has been made to utilize tokens in human vocabulary (i.e., tokens that are included in the LLM vocabulary). Previous work leverages items’ textual information such as titles (Bao et al., 2023), descriptions (Cui et al., 2022), and tags (Tan et al., 2024), aiming to utilize knowledge encoded in LLMs. More recently, utilizing external tokens for identifiers has attracted extensive attention due to its potential to include hierarchical information (Zhu et al., 2024; Zheng et al., 2024b; Wang et al., 2024e). To achieve this, existing work usually adopts hierarchical clustering or RQ-VAE (Lee et al., 2022) to obtain tokens in different granularity. Despite the effectiveness, token-sequence identifiers suffer from local optima issue and inference inefficiency. To improve efficiency, 2) single-token identifiers are proposed to represent each item with ID or semantic embedding (Li et al., 2023; Wang et al., 2024d). Nonetheless, existing work neither has poor generalization ability nor fails to capture CF information, thus leading to suboptimal results. In this work, we propose a novel set identifier paradigm, which employs a set of order-agnostic CF and semantic tokens. Two concurrent studies explore set identifiers for generative retrieval (Zeng et al., 2024; Zhang et al., 2024b), yet they still preserve token dependencies and heavily rely on autoregressive generation. Differently, our proposed paradigm achieves simultaneous generation without token dependencies, significantly enhancing generation efficiency.

### v2

@@ 1. Introduction
Token-sequence identifier utilizes a discrete token sequence to represent multi-dimensional item information. To generate items, LLMs use beam search to generate the top- item identifiers. Despite the effectiveness, token-sequence identifiers suffer from the 1) local optima issue (Zhou and Hansen, 2005) in the beam search. As illustrated in Figure 1, beam search greedily selects the sequence with top- probabilities at each generation step. However, the initial tokens of the target identifier might not necessarily align with the user preference. As such, the prefix of the target identifier has a low probability and will be pruned by beam search, causing inaccurate results. 2) Low generation efficiency in the autoregressive generation, which requires multiple serial LLM calls, thereby causing unaffordable computing burdens (Lin et al., 2024f) and severely hindering real-world deployments.
@@
To this end, we propose SETRec, an effective implementation of the set identifier paradigm. To integrate semantic and CF information, SETRec leverages CF and semantic tokenizers to assign each item with an order-agnostic token set containing CF and semantic embeddings. To eliminate token dependencies, 1) for user history encoding, we propose a special sparse attention mask, which discards the visibility of other tokens within identifiers and retains access to previous identifiers. 2) For simultaneous token generation, we introduce a query-guided generation mechanism, which adopts learnable vectors to guide LLMs to generate the embedding for each specific information dimension. 3) To ground the generated embedding set to existing items, SETRec collects embeddings from all items as grounding heads to obtain the item scores for ranking. We instantiate SETRec on T5 and Qwen and evaluate it on four real-world datasets under various scenarios (e.g., full ranking, warm- and cold-start ranking, and diverse item popularity groups) to demonstrate the effectiveness, efficiency, and generalization ability. Additionally, we evaluate SETRec on Qwen with different model sizes (i.e., 1.5B, 3B, and 7B), exhibiting promising scalability on cold-start items as model size increases. The code and datasets are available at https://github.com/Linxyhaha/SETRec.
@@ 2. Preliminaries
where is the -th token of identifier , represents the token sequence preceding , and is the LLM vocabulary. Despite the effectiveness, generating token sequences would result in the local optima issue and inference inefficiency. As shown in Figure 3, continuously increasing the beam size slightly improves recommendation accuracy, but remains inferior to globally optimal results. Worse still, the token-by-token generation requires multiple serial LLM calls, which significantly lowers the inference speed and hinders real-world applications.
@@ 4. Experiment
We carry out extensive experiments on four real-world datasets to answer the following research questions: RQ1: How does our proposed SETRec perform compared to different identifier baselines on different architectures of LLMs? RQ2: How do the different components of SETRec (i.e., CF embeddings, semantic embeddings, query vectors, and sparse attention) affect the performance? RQ3: How does SETRec perform when scaling up the model size and how does SETRec improve the overall performance? RQ4: How does SETRec perform with different number of semantic embeddings, tokenizer training strength, and semantic strength for inference?
@@ 4.1.1. Datasets
We conduct experiments on four real-world datasets across various domains. From Amazon review datasets222https://jmcauley.ucsd.edu/data/amazon/., we adopt three widely used benchmarks 1)Toys, 2) Beauty, and 3) Sports. The three Amazon datasets contain rich user interactions over a specific category of e-commerce products, where each item is associated with rich textual meta information such as title, description, category, and brand. In addition, we use a video games dataset 4) Steam333https://github.com/kang205/SASRec. proposed in (Kang and McAuley, 2018), which contains substantial user interactions on video games with abundant textual semantic information. For all datasets, we follow previous work (Wang et al., 2023a) to sort user interactions chronologically according to the timestamps and divide them into training, validation, and testing sets with a ratio of 8:1:1. In addition, we divide the items into warm and cold items444We denote warm- and cold-start items as warm and cold items for brevity., where the items that appear in the training set are warm items, otherwise cold items.
@@ 4.3.2. Item Group Analysis (RQ3)
To understand how SETRec improves performance, we evaluate it over items with different popularity. The performance comparison between SETRec and two competitive baselines from token-sequence identifiers (LETTER) and single-token identifiers (E4SRec) are reported in Figure 7. We can observe that 1) the performance gradually drops from G1 to G4. This makes sense since the less popular items have fewer interactions for LLMs to learn, thus leading to worse generation probabilities. Besides, 2) E4SRec outperforms LETTER on most popular items (G1) but usually yields inferior performance on unpopular items (G2-G4). This is due to that E4SRec only uses CF information, which relies on substantial interactions and therefore struggle on unpopular items. In contrast, LETTER additionally incorporates semantics into identifiers, thus achieving better generalization on sparse items. 3) SETRec consistently excels both E4SRec and LETTER over all groups. Notably, the improvements over sparse items are more significant, which partially explains the superiority of SETRec regarding overall performance.
@@ 4.3.4. Effect of Semantic Strength (RQ4)
To investigate how semantic information contributes to the performance during inference, we vary from to , where indicates that only CF score is used for ranking, and ranks items based solely on semantic scores (Eq. (7)). From the results in Figure 8, we can find that 1) incorporating semantic information during inference is necessary (inferior performance of than ), which facilitates global ranking over multi-dimensional information for strong generalization ability. Notably, 2) incorporating semantic scores brings more significant improvements on cold items, underscoring the semantic information for generalization. Moreover, 3) relying solely on semantics (), SETRec maintains competitive performance on warm items, which is attributed to the implicit alignment between CF and semantic tokens during training.
@@ 4.3.5. Hyper-parameter Sensitivity (RQ4)
1) Effect of . We vary the strength of AE loss for SETRec training and present the results on Toys in Figure 9(a). We can observe that the performance is overall improved when is increased from to , whereas significantly drops on warm items. Empirically, we recommend setting ranging from to . 2) Effect of . From the results in Figure 9(b), we can find that increasing semantic tokens generally improves the performance, possibly mitigating the potential information conflicts (Wang et al., 2024a) and embedding collapse issue (Guo et al., [n. d.]). However, blindly increasing the number of semantic tokens might hurt the performance. Because it is non-trivial to recover the category-level preference aligning well with the real-world scenarios (Lin et al., 2024b, e).
@@ 5. Related Work
LLM-based Recommendation. Harnessing LLMs for recommendations has garnered substantial attention across both academia and industry (Sun et al., 2024; Zhang et al., 2024c; Fu et al., 2024; Shi et al., 2024; Zhao et al., 2024; Zheng et al., 2024a; Li et al., 2024a; Gao et al., 2025; Li et al., 2024b; Xu et al., 2025; Wang et al., 2023b). Existing studies on LLM-based recommendations can be grouped into two research lines. 1) LLMs for discriminative recommendation, which typically aims to leverage LLMs to assist the conventional recommender models (Wang et al., 2024c; Ren et al., 2024a; Chen et al., 2024b; Wu et al., 2024). This line of work usually involves LLMs in different steps of recommendation pipelines (Lin et al., 2024a; Zhang et al., 2024a, b; Wang et al., 2024b) such as feature engineering (Xi et al., 2024; Ren et al., 2024b; Xi et al., 2024) and feature encoder (Chen et al., 2024a). 2) LLMs for generative recommendation, which regards LLMs as recommenders to directly generate items (Liao et al., 2024; Kim et al., 2024; Lin et al., 2024c; Wang et al., 2024f; Li et al., 2024a; Lin et al., 2024d). To build LLMs for generative recommendation, a key step is item tokenization, where each item is assigned an identifier for LLMs to encode user history and generate the next item. In this work, we critically analyze the fundamental principles of identifier design to achieve effective and efficient LLM-based generative recommendation.
@@
Identifier for LLM-based Recommendation. Existing identifier designs can be broadly categorized into two types: 1) token-sequence identifiers represent each item with a discrete token sequence. Early studies leverage items’ textual information such as titles (Bao et al., 2023), descriptions (Cui et al., 2022), and tags (Tan et al., 2024), aiming to utilize knowledge encoded in LLMs through human vocabulary. More recently, utilizing external tokens for identifiers has attracted extensive attention due to its potential to include hierarchical information (Zhu et al., 2024; Zheng et al., 2024b; Wang et al., 2024e). Despite the effectiveness and the potential to incorporate both semantic and CF information (Zhang et al., 2025; Wang et al., 2024a), token-sequence identifiers suffer from local optima issue and inference inefficiency. To improve efficiency, 2) single-token identifiers are proposed to represent each item with ID or semantic embedding (Li et al., 2023; Wang et al., 2024d). Nonetheless, existing work neither has poor generalization ability nor fails to capture CF information. In this work, we propose a novel set identifier paradigm, which employs a set of order-agnostic CF and semantic tokens. Two concurrent studies explore set identifiers for generative retrieval (Zeng et al., 2024; Zhang et al., 2024b), yet they still preserve token dependencies and heavily rely on autoregressive generation. Differently, our proposed paradigm achieves simultaneous generation without token dependencies, significantly enhancing generation efficiency.